paul is rebuilding his mailbox post. The horizontal portion is 14inches, where a is 10 inches. The vertical post is 49 inches, Where b is 13inches. Those two pieces meet at a right angle. What is the minimum lenght necessary for the angled support, c ?
1
Expert's answer
2011-11-01T11:03:19-0400
minimum lenght c=sqrt(10*10+13*13)=sqrt(100+169)=sqrt(269).